# Planting and MCMC Sampling from the Potts model

### v1

@@ Appendix A Remaining Proofs for Section 2
    $\beta<\beta_{u}^{\prime}$: if $\mathbb{P}\big{[}\big{(}\hat{\boldsymbol{G}}(\hat{\boldsymbol{\sigma}}_{\mathrm{d}}),\hat{\boldsymbol{\sigma}}_{\mathrm{d}}\big{)}\in{\mathcal{E}}\big{]}\leq g(n)$, then, whp over $\boldsymbol{G}$, $\mathbb{P}\big{[}(\boldsymbol{G},\boldsymbol{\sigma}_{\boldsymbol{G},\mathrm{d}}\big{)}\in{\mathcal{E}}\mid\boldsymbol{G}\big{]}\leq\frac{g(n)}{f(n)}$.
@@
    $\beta>\beta_{u}$: if $\mathbb{P}\big{[}\big{(}\hat{\boldsymbol{G}}(\hat{\boldsymbol{\sigma}}_{\mathrm{o}}),\hat{\boldsymbol{\sigma}}_{\mathrm{o}}\big{)}\in{\mathcal{E}}\big{]}\leq g(n)$, then, whp over $\boldsymbol{G}$, $\mathbb{P}\big{[}(\boldsymbol{G},\boldsymbol{\sigma}_{\boldsymbol{G},\mathrm{o}}\big{)}\in{\mathcal{E}}\mid\boldsymbol{G}\big{]}\leq\frac{g(n)}{f(n)}$.
@@
We give the proof for the regime $\beta<\beta_{u}^{\prime}$, the proof for $\beta>\beta_{u}$ is completely analogous. Let $\mathcal{Q}$ be the set consisting of $G\in\mathcal{G}_{n,d}$ that satisfy
@@ Appendix B Remaining Proofs for Section 3
We prove the first equation, the second is analogous. Fix $\beta<\beta_{u}^{\prime}$. By Lemma 2.3, for the disordered planted model $\mathbb{P}\big{[}\big{\|}\rho^{\hat{\boldsymbol{G}}(\hat{\boldsymbol{\sigma}}_{\mathrm{d}}),\hat{\boldsymbol{\sigma}}_{\mathrm{d}}}-\rho_{\mathrm{d}}\big{\|}_{1}=o(1)\big{]}=1-\mathrm{e}^{-\Omega(n)}$. So Lemma 2.2 yields that $\mathbb{P}\Big{[}\big{\|}\rho^{\boldsymbol{G},\boldsymbol{\sigma}_{\boldsymbol{G},\mathrm{d}}}-\rho_{\mathrm{d}}\big{\|}_{1}=o(1)\,\Big{|}\,\boldsymbol{G}\Big{]}=1-\mathrm{e}^{-\Omega(n)}$ whp over $\boldsymbol{G}$. Therefore, for such $\boldsymbol{G}$, conditioned on $\boldsymbol{\sigma}\in S_{\mathrm{d}}$, the number of monochromatic edges under $\boldsymbol{\sigma}$ is at most $n\big{(}\tfrac{d}{2}\sum_{i\in[q]}\rho_{\mathrm{d},ii}\big{)}+n\rho/4=n\frac{1}{p}\mathfrak{m}_{\mathrm{d}}(\beta)+n\rho/4$ w.p. $1-\mathrm{e}^{-\Omega(n)}$. Therefore after the percolation step, by standard Chernoff bounds, we have $|\boldsymbol{F}|\leq n\mathfrak{m}_{\mathrm{d}}(\beta)+n\rho/4$ w.p. $1-\mathrm{e}^{-\Omega(n)}$. Since $\mathfrak{m}_{\mathrm{d}}(\beta)\leq\mathfrak{m}_{\mathrm{d}}(\beta_{c})$, by the Definition 3.1 of $\Omega_{\boldsymbol{G},\mathrm{d}},\Omega_{\boldsymbol{G},\mathrm{o}}$ we conclude that $\mathbb{P}[\boldsymbol{F}\in\Omega_{\boldsymbol{G},\mathrm{d}}\mid\boldsymbol{\sigma}\in S_{\mathrm{d}}]=1-\mathrm{e}^{-\Omega(n)}$ and $\mathbb{P}[\boldsymbol{F}\in\Omega_{\boldsymbol{G},\mathrm{o}}\mid\boldsymbol{\sigma}\in S_{\mathrm{d}}]=\mathrm{e}^{-\Omega(n)}$. ∎
@@
If $\beta\leq\beta_{c}$, for $\boldsymbol{F}_{\mathrm{d}}\sim\pi_{G,\mathrm{d}}$, it holds that $\mathbb{P}(|\boldsymbol{F}_{\mathrm{d}}|\geq n\mathfrak{m}_{\mathrm{d}}+n\delta)=\mathrm{e}^{-\Omega(n)}$; if $\beta\geq\beta_{c}$, for $\boldsymbol{F}_{\mathrm{o}}\sim\pi_{G,\mathrm{o}}$, it holds that $\mathbb{P}(|\boldsymbol{F}_{\mathrm{o}}|\leq n\mathfrak{m}_{\mathrm{d}}-n\delta)=\mathrm{e}^{-\Omega(n)}$.
@@ Appendix C Remaining Proofs for Section 4
First of all, it is well-known that whp $G\sim\mathcal{G}_{n,d}$ is locally treelike (see e.g. [BG, Fact 2.3]), i.e. that there exists a constant $L=L(\delta,d)$ such that for all $v$, we can remove at most $L$ edges from $B_{\ell}(v)$ to make it a tree. Second, by Lemma 4.2, there exists $K=K(d,q,\beta,\delta)$ such that whp over $G\sim\mathcal{G}_{n,d}$, with probability $\geq 1-n^{-3}$, for all $v$, a random disordered configuration is $K$-shattered at distance $\ell$ from $v$. Thus from now on we assume $G$ satisfies these two properties.
@@
For the case $\beta<\beta_{u}$, this bound was proven in [BG, Proposition 3.3], however parts of their proof use estimates that only hold in uniqueness. We are able to extend their bound for all $\beta$ by closely following the start of their argument and then using different estimates which continue to hold beyond the uniqueness threshold.
The starting point for the proof is two results from [BG]: Lemma 5.3 and equation (5.4). Lemma 5.3 bounds the TV-distance on marginals on $e$ by $\pi_{G,v,\ell}^{\eta}(\Upsilon_{G,v,\ell}^{\eta})$, where $\Upsilon_{G,v,\ell}^{\eta}$ is the event such that a configuration $\boldsymbol{F}$ satisfies the following: for all $1\leq i\leq\ell$, there are two distinct vertices $u_{1},u_{2}\in S_{i}(v)$ and two corresponding paths $\gamma_{i,1},\gamma_{i,2}$ such that for both $j\in\{1,2\}$:
@@
Then in the proof of [BG, Proposition 3.3], $\pi_{G,v,\ell}^{\eta}(\Upsilon_{G,v,\ell}^{\eta})$ is further bound by a probability of an event roughly corresponding to two disjoint paths from a neighbour of $v$ to $S_{\ell}(v)$. In order to define this event, we need to introduce more notation.
Since $B_{\ell}(v)$ is $L$-treelike, there is a set $E_{\mathcal{E}}\subseteq E(B_{\ell}(v))$ of at most $L$ edges such that removing $E_{\mathcal{E}}$ makes the ball a tree. Thus there are also $k\leq 2L$ distances $1\leq d_{1}<\dots<d_{k}\leq\ell$ such that $S_{d_{i}}(v)$ contains a vertex from $V(E_{\mathcal{E}})$. Also denote $d_{0}=1,d_{k+1}=\ell$. The following bound is shown in [BG, Equation (5.4)]:
@@
Since $\hat{p}^{-8L}$ is a constant, as a final step it suffices to show that $\frac{\pi^{\eta}_{G,v,\ell}(\Gamma\cup\Gamma^{\prime}\in\cdot)}{\tilde{\pi}(\Gamma\cup\Gamma^{\prime}\in\cdot)}$ is bounded by a constant (not depending on $(\Gamma;\Gamma^{\prime})$). Consider any configuration $A\subseteq E(B_{\ell}(v))$. Since $|E_{\mathcal{E}}|\leq L$, $\eta$ is $K$-shattered, and adding an edge can decrease the number of components by at most one, we get that $|A\setminus E_{\mathcal{E}}|\leq|A|\leq|A\setminus E_{\mathcal{E}}|+L$, and the number of components in the ball for configuration $A$ w.r.t. boundary condition $\eta$ differ by at most $K+L$ from the number of components for configuration $A\setminus E_{\mathcal{E}}$ w.r.t. to the free boundary condition. Thus the weights of $A$ w.r.t. $\eta$ and of $A\setminus E_{\mathcal{E}}$ w.r.t. to -- differ by a factor of at most $q^{L+K}[\max(\mathrm{e}^{\beta}-1,(\mathrm{e}^{\beta}-1)^{-1})]^{L}$.
Given that there are at most $2^{|E_{\mathcal{E}}|}\leq 2^{L}$ configurations $A\subseteq E(B_{\ell}(v))$ agreeing on $A\setminus E_{\mathcal{E}}$, we get that the cumulative weights of the event “$\Gamma\cup\Gamma^{\prime}$ are all in” differ by a constant factor (independent of $(\Gamma;\Gamma^{\prime})$) when w.r.t. $\eta$, and w.r.t. -- when restricted to all edges in $E_{\mathcal{E}}$ being out, and so do their respective partition functions. It hence follows that $\frac{\pi^{\eta}_{G,v,\ell}(\Gamma\cup\Gamma^{\prime}\in\cdot)}{\tilde{\pi}(\Gamma\cup\Gamma^{\prime}\in\cdot)}$ is bounded by a constant, which concludes the proof. ∎
@@
In this section we give the proof of Lemma 5.1. We first consider the case where $\boldsymbol{d}=(d_{1},\ldots,d_{n})$ is the degree sequence of a $d$-regular graph (i.e., $d_{i}=d$ for all $i$). For simplicity, in this uniform setting we denote the percolation model with $m=pdn/2+o(n)$ edges by $\tilde{\boldsymbol{G}}_{p}$. The following bounds were shown in [coja2023].
Lemma D.1 ([coja2023, Proposition 5.4]).
@@
[…]
@@
Observe that the graph $\tilde{\boldsymbol{G}}_{n,\boldsymbol{d},m}$, conditional on the number of percolated edges from vertices of $S$ being (some) integer $k$, the graph $\tilde{\boldsymbol{G}}_{n,\boldsymbol{d},m}\setminus S$ is distributed as an exact ($d$-regular) edge model with $n-|S|$ vertices and $m-k$ edges – as every set of $m-k$ edges (submatchings) is equally likely by symmetry. Since $|S|=O(n^{1/6})$, we trivially have that $k=O(n^{1/6})$ as well. Thus, for any such $k$, it holds that $\frac{m-k}{dn/2}=p+o(1)-O(n^{-5/6})>\frac{1}{d-1}$ for sufficiently large $n$. Furthermore, since $\chi(p)$ is continuous on $(\tfrac{1}{d-1},1)$, it also holds that $\chi^{\prime}:=\chi(\frac{m-k}{dn/2})=\chi(p)\pm\delta/3$. Applying therefore Lemma D.1 to $\tilde{\boldsymbol{G}}_{n,\boldsymbol{d},m}\setminus S$, we get that with probability $1-\mathrm{e}^{-\Omega(n)}$, its largest component has size $(\chi^{\prime}\pm\delta/3)(n-|S|)=(\chi\pm\delta)n$ for $n$ large enough, and hence so does $\tilde{\boldsymbol{G}}_{n,\boldsymbol{d},m}$.
@@
Let $0<d_{0}\leq d$ be integers, and $p\in(\tfrac{1}{d-1})$ be a real. Then consider a particular case of Galton-Watson process where the root has offspring distribution ${\rm Bin}(d_{0},p)$, and in all the following generations have offspring distribution ${\rm Bin}(d,p)$. Let $Z_{r}$ be the size of the surviving population at generation $r$. Then, for any fixed integer $K>0$, $\mathbb{P}(Z_{r}\geq K\mid Z_{r}>0)=1-o_{r}(1)$.
@@
Note that once at second or later generation, the process is supercritical, so then we can apply the Lemma D.3. Thus we need to consider the case of the root.
Note that if $Z_{r}>0$, then there is a non-empty set ${\mathcal{C}}$ of children of the root whose branching process survive $r-1$ generations. Since their individual branching processes – denote them $(Z_{t}^{i})_{i\in{\mathcal{C}},t\geq 0}$, are independent, $\mathbb{P}(Z_{r}<K\mid Z_{r}>0,\mathcal{S}\text{ survive})\leq\prod_{i\in{\mathcal{C}}}\mathbb{P}(Z_{r-1}^{i}<K\mid Z_{r-1}^{i}>0)=o_{r}(1)$ by the previous lemma. Summing over the possible (non-empty) sets of surviving children, we obtain that the bound holds unconditionally. ∎
@@
    If $p<1/(d-1)$, $\mathbb{P}\big{[}{\mathcal{C}}_{1}(\tilde{\boldsymbol{G}}_{n,\boldsymbol{d},m})\geq Mn^{1/2}]\leq\mathrm{e}^{-\Omega(n)}$.
@@
    If $p>1/(d-1)$, $\mathbb{P}\big{[}{\mathcal{C}}_{1}(\tilde{\boldsymbol{G}}_{n,\boldsymbol{d},m})\geq(\chi-\delta)n]\geq 1-\mathrm{e}^{-\Omega(n)}$ and $\mathbb{P}\big{[}{\mathcal{C}}_{2}(\tilde{\boldsymbol{G}}_{n,\boldsymbol{d},m})\geq Mn^{2/3}]\leq\mathrm{e}^{-\Omega(n)}$. Moreover, with $\varphi=\varphi(p)$ as in (5.1), for any subset $S^{\prime}\subseteq S$,
@@
For Item (1) (the subcritical case), the statement is proved in [coja2023, Proposition 5.2] for $S=\emptyset$. Note that while the statement of the proposition is for $d$-regular degree sequences, as it is noted therein, it also applies for percolation on a graph of maximum degree $d$ (with any degree sequence).
We thus focus on Item (2). The first part of the claim follows from Lemma D.2. It only remains to bound the probability that none of the vertices in $S^{\prime}$ belong to the largest component. In the proof, we work with a number of sufficiently small, or sufficiently large (in terms of $d,p,\epsilon$ constants) constants, whose values are specified later. Namely, $R,K>0$ are sufficiently large constants, and $\epsilon_{0},\epsilon_{1}>0$ are sufficiently small.
@@
The proof of the claim is standard and similar to the proof of [Blanca2, Lemma 15] so we sketch the argument briefly. Consider a BFS revealing process from all the vertices in $S^{\prime}$ up to depth $R$. Say that a “bad event” is revealing a match going to a vertex that has been either already seen or belongs to $S$ (only vertices in $S$ can have degree $\neq d$). Note that the size of each neighbourhood is at most $|S^{\prime}|d^{R}$ (at least for $R$ large enough), and thus in each revealing step, regardless of previous outcomes, the probability of a bad event occurring is bounded by $\frac{|S^{\prime}|d^{R}+|S|}{n-|S^{\prime}|d^{R}-|S|}\leq\frac{2|S|d^{R}}{n}$ (since $S=O(n^{1/6})$, for large enough $n$ the denominator is at least $n/2$). Hence we can upper bound the number of bad events by a sum of $|S^{\prime}|d^{R}$ i.i.d. Bernoulli random variables with mean $\frac{2|S|d^{R}}{n}$. Then by Chernoff bounds (e.g. [Blanca2, Lemma 16]), (#bad $\mathbb{P}(\#\text{bad events}\geq\epsilon_{0}|S^{\prime}|/2)\leq\exp\big{\{}-\frac{\epsilon_{0}|S^{\prime}|}{2}\log\big{(}\frac{n}{2e|S|d^{R}}\big{)}\big{\}}$. Since $2e|S|d^{R}=O(n^{1/6})$, the probability is at most $(C^{\prime}n^{-5\epsilon_{0}/12})^{|S^{\prime}|}$ for some constant $C^{\prime}>0$ (independent of $n$). Note that each bad event makes at most two neighbourhoods not nice, and that this probability is at most $n^{-C^{\prime}|S^{\prime}|}$ for some constant $C^{\prime}(C,\epsilon_{0})>0$. Thus with probability at least $1-n^{-C|S^{\prime}|}$ there are at least $(1-\epsilon_{0})|S^{\prime}|$ nice neighbourhoods, which concludes the claim.
@@
Each of surviving leaves has $d-1$ endpoints with an unrevealed match. Notice that all of these $L_{v}$ leaves are connected in $\tilde{\boldsymbol{G}}_{n,\boldsymbol{d},m}$, and thus, crucially, all of their endpoints can be seen as connected, or belonging to the same vertex. We can hence “regroup” the endpoints for the surviving leaves corresponding to each $v\in T$ into groups of $d$ endpoints, thus obtaining $\lfloor\frac{L_{v}(d-1)}{d}\rfloor$ new vertices of degree $d$ (we can “discard” the excess endpoints). These new vertices will be the vertices in $\boldsymbol{H}$ corresponding to $v$.

### v2

@@ Appendix A Remaining Proofs for Section 2
    $\beta\leq\beta_{c}$: if $\mathbb{P}\big[\big(\hat{\boldsymbol{G}}(\hat{\boldsymbol{\sigma}}_{\mathrm{d}}),\hat{\boldsymbol{\sigma}}_{\mathrm{d}}\big)\in{\mathcal{E}}\big]\leq g(n)$, then, whp over $\boldsymbol{G}$, $\mathbb{P}\big[(\boldsymbol{G},\boldsymbol{\sigma}_{\boldsymbol{G},\mathrm{d}}\big)\in{\mathcal{E}}\mid\boldsymbol{G}\big]\leq\frac{g(n)}{f(n)}$.
@@
    $\beta\geq\beta_{c}$: if $\mathbb{P}\big[\big(\hat{\boldsymbol{G}}(\hat{\boldsymbol{\sigma}}_{\mathrm{o}}),\hat{\boldsymbol{\sigma}}_{\mathrm{o}}\big)\in{\mathcal{E}}\big]\leq g(n)$, then, whp over $\boldsymbol{G}$, $\mathbb{P}\big[(\boldsymbol{G},\boldsymbol{\sigma}_{\boldsymbol{G},\mathrm{o}}\big)\in{\mathcal{E}}\mid\boldsymbol{G}\big]\leq\frac{g(n)}{f(n)}$.
@@
We give the proof for the first item, the proof for the other is completely analogous. Fix arbitrary positive $\beta\leq\beta_{c}$ and let $\mathcal{Q}$ be the set consisting of $G\in\mathcal{G}_{n,d}$ that satisfy
@@ Appendix B Remaining Proofs for Section 3
We prove the first equation, the second is analogous. Fix $\beta\leq\beta_{c}$. By Lemma 2.3, for the disordered planted model $\mathbb{P}\big[\big\|\rho^{\hat{\boldsymbol{G}}(\hat{\boldsymbol{\sigma}}_{\mathrm{d}}),\hat{\boldsymbol{\sigma}}_{\mathrm{d}}}-\rho_{\mathrm{d}}\big\|_{1}=o(1)\big]=1-\mathrm{e}^{-\Omega(n)}$. So Lemma 2.2 yields that $\mathbb{P}\Big[\big\|\rho^{\boldsymbol{G},\boldsymbol{\sigma}_{\boldsymbol{G},\mathrm{d}}}-\rho_{\mathrm{d}}\big\|_{1}=o(1)\,\Big|\,\boldsymbol{G}\Big]=1-\mathrm{e}^{-\Omega(n)}$ whp over $\boldsymbol{G}$. Therefore, for such $\boldsymbol{G}$, conditioned on $\boldsymbol{\sigma}\in S_{\mathrm{d}}$, the number of monochromatic edges under $\boldsymbol{\sigma}$ is at most $n\big(\tfrac{d}{2}\sum_{i\in[q]}\rho_{\mathrm{d},ii}\big)+n\rho/4=n\frac{1}{p}\mathfrak{m}_{\mathrm{d}}(\beta)+n\rho/4$ w.p. $1-\mathrm{e}^{-\Omega(n)}$. Therefore after the percolation step, by standard Chernoff bounds, we have $|\boldsymbol{F}|\leq n\mathfrak{m}_{\mathrm{d}}(\beta)+n\rho/4$ w.p. $1-\mathrm{e}^{-\Omega(n)}$. Since $\mathfrak{m}_{\mathrm{d}}(\beta)\leq\mathfrak{m}_{\mathrm{d}}(\beta_{c})$, by the Definition 3.1 of $\Omega_{\boldsymbol{G},\mathrm{d}},\Omega_{\boldsymbol{G},\mathrm{o}}$ we conclude that $\mathbb{P}[\boldsymbol{F}\in\Omega_{\boldsymbol{G},\mathrm{d}}\mid\boldsymbol{\sigma}\in S_{\mathrm{d}}]=1-\mathrm{e}^{-\Omega(n)}$ and $\mathbb{P}[\boldsymbol{F}\in\Omega_{\boldsymbol{G},\mathrm{o}}\mid\boldsymbol{\sigma}\in S_{\mathrm{d}}]=\mathrm{e}^{-\Omega(n)}$. ∎
@@
If $\beta\leq\beta_{c}$, for $\boldsymbol{F}_{\mathrm{d}}\sim\pi_{G,\mathrm{d}}$, it holds that $\mathbb{P}(|\boldsymbol{F}_{\mathrm{d}}|\geq n\mathfrak{m}_{\mathrm{d}}+n\delta)=\mathrm{e}^{-\Omega(n)}$; if $\beta\geq\beta_{c}$, for $\boldsymbol{F}_{\mathrm{o}}\sim\pi_{G,\mathrm{o}}$, it holds that $\mathbb{P}(|\boldsymbol{F}_{\mathrm{o}}|\leq n\mathfrak{m}_{\mathrm{o}}-n\delta)=\mathrm{e}^{-\Omega(n)}$.
@@ Appendix C Remaining Proofs for Section 4
In this section we prove Lemma 4.3. Throughout this proof, a nontrivial boundary component of $\eta$ (where $\eta$ is a boundary condition induced by $F\setminus B_{\ell}(v)$ for some $F\subseteq E$) means a a component of $(V,F\backslash E(B_{\ell}(v)))$ containing at least two vertices of $S_{\ell}(v)$.
@@
First of all, it is well-known that whp $G\sim\mathcal{G}_{n,d}$ is locally treelike (see e.g. [9, Fact 2.3]), i.e. that there exists a constant $L=L(\delta,d)$ such that for all $v$, we can remove at most $L$ edges from $B_{\ell}(v)$ to make it a tree. Second, by Lemma 4.2, there exists $K=K(d,q,\beta,\delta)$ such that whp over $G\sim\mathcal{G}_{n,d}$, with probability $\geq 1-n^{-3}$, for all $v$, a random disordered configuration is $K$-shattered at distance $\ell$ from $v$. Thus from now on we assume $G$ satisfies these two properties.
@@
For the case $\beta<\beta_{u}$, this bound was proven in [9, Proposition 3.3], however parts of their proof use estimates that only hold in uniqueness. We are able to extend their bound for all $\beta$ by closely following the start of their argument and then using different estimates which continue to hold beyond the uniqueness threshold.
The starting point for the proof is two results from [9]: Lemma 5.3 and equation (5.4). Lemma 5.3 bounds the TV-distance on marginals on $e$ by $\pi_{G,v,\ell}^{\eta}(\Upsilon_{G,v,\ell}^{\eta})$, where $\Upsilon_{G,v,\ell}^{\eta}$ is the event such that a configuration $\boldsymbol{F}$ satisfies the following: for all $1\leq i\leq\ell$, there are two distinct vertices $u_{1},u_{2}\in S_{i}(v)$ and two corresponding paths $\gamma_{i,1},\gamma_{i,2}$ such that for both $j\in\{1,2\}$:
@@
Then in the proof of [9, Proposition 3.3], $\pi_{G,v,\ell}^{\eta}(\Upsilon_{G,v,\ell}^{\eta})$ is further bound by a probability of an event roughly corresponding to two disjoint paths from a neighbour of $v$ to $S_{\ell}(v)$. In order to define this event, we need to introduce more notation.
Since $B_{\ell}(v)$ is $L$-treelike, there is a set $E_{\mathcal{E}}\subseteq E(B_{\ell}(v))$ of at most $L$ edges such that removing $E_{\mathcal{E}}$ makes the ball a tree. Thus there are also $k\leq 2L$ distances $1\leq d_{1}<\dots<d_{k}\leq\ell$ such that $S_{d_{i}}(v)$ contains a vertex from $V(E_{\mathcal{E}})$. Also denote $d_{0}=1,d_{k+1}=\ell$. The following bound is shown in [9, Equation (5.4)]:
@@
Since $\hat{p}^{-8L}$ is a constant, as a final step it suffices to show that $\frac{\pi^{\eta}_{G,v,\ell}(\Gamma\cup\Gamma^{\prime}\in\cdot)}{\tilde{\pi}(\Gamma\cup\Gamma^{\prime}\in\cdot)}$ is bounded by a constant (not depending on $(\Gamma;\Gamma^{\prime})$). Consider any configuration $A\subseteq E(B_{\ell}(v))$. Since $|E_{\mathcal{E}}|\leq L$, $\eta$ is $K$-shattered, and adding an edge can decrease the number of components by at most one, we get that $|A\setminus E_{\mathcal{E}}|\leq|A|\leq|A\setminus E_{\mathcal{E}}|+L$, and the number of components in the ball for configuration $A$ w.r.t. boundary condition $\eta$ differ by at most $K+L$ from the number of components for configuration $A\setminus E_{\mathcal{E}}$ w.r.t. to the free boundary condition. Thus the weights of $A$ w.r.t. $\eta$ and of $A\setminus E_{\mathcal{E}}$ w.r.t. to - differ by a factor of at most $q^{L+K}[\max(\mathrm{e}^{\beta}-1,(\mathrm{e}^{\beta}-1)^{-1})]^{L}$.
Given that there are at most $2^{|E_{\mathcal{E}}|}\leq 2^{L}$ configurations $A\subseteq E(B_{\ell}(v))$ agreeing on $A\setminus E_{\mathcal{E}}$, we get that the cumulative weights of the event $\Gamma\cup\Gamma^{\prime}$ are all in” differ by a constant factor (independent of $(\Gamma;\Gamma^{\prime})$) when w.r.t. $\eta$, and w.r.t. - when restricted to all edges in $E_{\mathcal{E}}$ being out, and so do their respective partition functions. It hence follows that $\frac{\pi^{\eta}_{G,v,\ell}(\Gamma\cup\Gamma^{\prime}\in\cdot)}{\tilde{\pi}(\Gamma\cup\Gamma^{\prime}\in\cdot)}$ is bounded by a constant, which concludes the proof. ∎
@@
In this section we give the proof of Lemma 5.1. We first consider the case where $\boldsymbol{d}=(d_{1},\ldots,d_{n})$ is the degree sequence of a $d$-regular graph (i.e., $d_{i}=d$ for all $i$). For simplicity, in this uniform setting we denote the percolation model with $m=pdn/2+o(n)$ edges by $\tilde{\boldsymbol{G}}_{p}$. The following bounds were shown in [14].
Lemma D.1 ([14, Proposition 5.4]).
@@
$\mathbb{P}\big(\mbox{$|{\mathcal{C}}_{1}(\tilde{\boldsymbol{G}}_{n,\boldsymbol{d},m})|\geq(\chi\pm\delta)n$ and $|{\mathcal{C}}_{2}(\tilde{\boldsymbol{G}}_{n,\boldsymbol{d},m})|\leq Mn^{2/3}$}\big)\geq 1-\mathrm{e}^{-\Omega(n)}.$
@@
Observe that the graph $\tilde{\boldsymbol{G}}_{n,\boldsymbol{d},m}$, conditional on the number of percolated edges from vertices of $S$ being (some) integer $k$, the graph $\tilde{\boldsymbol{G}}_{n,\boldsymbol{d},m}\setminus S$ is distributed as an exact ($d$-regular) edge model with n|S|n-|S| vertices and mkm-k edges – as every set of mkm-k edges (submatchings) is equally likely by symmetry. Since $|S|=O(n^{1/6})$, we trivially have that $k=O(n^{1/6})$ as well. Thus, for any such $k$, it holds that $\frac{m-k}{dn/2}=p+o(1)-O(n^{-5/6})>\frac{1}{d-1}$ for sufficiently large $n$. Furthermore, since $\chi(p)$ is continuous on $(\tfrac{1}{d-1},1)$, it also holds that $\chi^{\prime}:=\chi(\frac{m-k}{dn/2})=\chi(p)\pm\delta/3$. Applying therefore Lemma D.1 to $\tilde{\boldsymbol{G}}_{n,\boldsymbol{d},m}\setminus S$, we get that with probability $1-\mathrm{e}^{-\Omega(n)}$, its largest component has size $(\chi^{\prime}\pm\delta/3)(n-|S|)=(\chi\pm\delta)n$ for $n$ large enough, and hence so does $\tilde{\boldsymbol{G}}_{n,\boldsymbol{d},m}$.
@@
Let $0<d_{0}\leq d$ be integers, and $p\in(\tfrac{1}{d-1})$ be a real. Then consider a Galton-Watson process where the root has offspring distribution ${\rm Bin}(d_{0},p)$, and all the following generations have offspring distribution ${\rm Bin}(d,p)$. Let $Z_{r}$ be the size of the surviving population at generation $r$. Then, for any fixed integer $K>0$, $\mathbb{P}(Z_{r}\geq K\mid Z_{r}>0)=1-o_{r}(1)$.
@@
Note that once at second or later generation, the process is supercritical, so then we can apply Lemma D.3. Thus we need to consider the case of the root.
Note that if $Z_{r}>0$, then there is a non-empty set $\mathcal{C}$} of children of the root whose branching process survive r1r-1 generations. Since their individual branching processes – denote them $(Z_{t}^{i})_{i\in{\mathcal{C}},t\geq 0}$, are independent, $\mathbb{P}(Z_{r}<K\mid Z_{r}>0,\mathcal{S}\text{ survive})\leq\prod_{i\in{\mathcal{C}}}\mathbb{P}(Z_{r-1}^{i}<K\mid Z_{r-1}^{i}>0)=o_{r}(1)$ by the previous lemma. Summing over the possible (non-empty) sets of surviving children, we obtain that the bound holds unconditionally. ∎
@@
    If $p<\tfrac{1}{d-1}$, $\mathbb{P}\big[{\mathcal{C}}_{1}(\tilde{\boldsymbol{G}}_{n,\boldsymbol{d},m})\geq Mn^{1/2}]\leq\mathrm{e}^{-\Omega(n)}$.
@@
    If $p>\frac{1}{d-1}$, $\mathbb{P}\big[{\mathcal{C}}_{1}(\tilde{\boldsymbol{G}}_{n,\boldsymbol{d},m})\leq(\chi-\delta)n]\leq\mathrm{e}^{-\Omega(n)}$ and $\mathbb{P}\big[{\mathcal{C}}_{2}(\tilde{\boldsymbol{G}}_{n,\boldsymbol{d},m})\geq Mn^{2/3}]\leq\mathrm{e}^{-\Omega(n)}$. Moreover, with $\varphi=\varphi(p)$ as in (5.1), for any subset $S^{\prime}\subseteq S$,
@@
For Item (1) (the subcritical case), the statement is proved in [14, Proposition 5.2] for $S=\emptyset$. Note that while the statement of the proposition is for $d$-regular degree sequences, as it is noted therein, it also applies for percolation on a graph of maximum degree $d$ (with any degree sequence).
We thus focus on Item (2). The first part of the claim follows from Lemma D.2. It only remains to bound the probability that none of the vertices in $S^{\prime}$ belong to the largest component. Let $R,K>0$ be sufficiently large constants and $\epsilon_{0},\epsilon_{1}>0$ be sufficiently small, to be specified later.
@@
The proof of the claim is standard and similar to the proof of [8, Lemma 15] so we sketch the argument briefly. Consider a BFS revealing process from all the vertices in $S^{\prime}$ up to depth $R$. Say that a “bad event” is revealing a match going to a vertex that has been either already seen or belongs to $S$ (only vertices in $S$ can have degree $\neq d$). Note that the size of each neighbourhood is at most $|S^{\prime}|d^{R}$ (at least for $R$ large enough), and thus in each revealing step, regardless of previous outcomes, the probability of a bad event occurring is bounded by $\frac{|S^{\prime}|d^{R}+|S|}{n-|S^{\prime}|d^{R}-|S|}\leq\frac{2|S|d^{R}}{n}$ (since $S=O(n^{1/6})$, for large enough $n$ the denominator is at least $n/2$). Hence we can upper bound the number of bad events by a sum of $|S^{\prime}|d^{R}$ i.i.d. Bernoulli random variables with mean $\frac{2|S|d^{R}}{n}$. Then by Chernoff bounds (e.g. [8, Lemma 16]), $\mathbb{P}(\#\text{bad events}\geq\epsilon_{0}|S^{\prime}|/2)\leq\exp\big\{-\frac{\epsilon_{0}|S^{\prime}|}{2}\log\big(\frac{n}{2e|S|d^{R}}\big)\big\}$. Since $2e|S|d^{R}=O(n^{1/6})$, the probability is at most $(C^{\prime}n^{-5\epsilon_{0}/12})^{|S^{\prime}|}$ for some constant $C^{\prime}>0$ (independent of $n$). Note that each bad event makes at most two neighbourhoods not nice, and that this probability is at most $n^{-C^{\prime}|S^{\prime}|}$ for some constant $C^{\prime}(C,\epsilon_{0})>0$. Thus with probability at least $1-n^{-C|S^{\prime}|}$ there are at least $(1-\epsilon_{0})|S^{\prime}|$ nice neighbourhoods, which concludes the claim.
@@
Each of surviving leaves has d1d-1 endpoints with an unrevealed match. Notice that all of these $L_{v}$ leaves are connected in $\tilde{\boldsymbol{G}}_{n,\boldsymbol{d},m}$, and thus, crucially, all of their endpoints can be seen as connected, or belonging to the same vertex. We can hence “regroup” the endpoints for the surviving leaves corresponding to each $v\in T$ into groups of $d$ endpoints, thus obtaining $\lfloor\frac{L_{v}(d-1)}{d}\rfloor$ new vertices of degree $d$ (we can “discard” the excess endpoints). These new vertices will be the vertices in $\boldsymbol{H}$ corresponding to $v$.